OpenStudy (anonymous):

What are the first five terms of the sequence given by the formula an = 5n + 1? (1 point) 1, 5, 9, 13, 17 6, 11, 16, 21, 26 5, 10, 15, 20, 25 5, 6, 7, 8, 9

OpenStudy (anonymous):

\[an = 5n + 1\] \[a1 = 5(1) + 1\] \[a1 = 5 + 1\] \[a1 = 6\]

OpenStudy (anonymous):

@some_someone thanks so much.

OpenStudy (anonymous):

\[an = 5n + 1\] \[a2 = 5(2) + 1\] \[a2 = 10 + 1\] \[a2 = 11\]

OpenStudy (anonymous):

\[an = 5n + 1\] \[a3 = 5(3) + 1\] \[a3 = 15 + 1\] \[a3 = 16\]

OpenStudy (anonymous):

\[an = 5n + 1\] \[a4 = 5(4) + 1\] \[a4 = 20 + 1\] \[a4 = 21\]

OpenStudy (anonymous):

\[an = 5n + 1\] \[a5 = 5(5) + 1\] \[a5 = 25 + 1\] \[a5 = 26\]